Grade 11Inorganic Chemistry

If the equilibrium constant for A+2B + 5/2D is 4.0 then what is the value of the equilibrium constant for the reaction 2C+ 5D 2A + 4B at the same temperature?

kiara
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
A + 2B → C + (5/2)D …… K₁ = [C] [D]^(5/2) / {[A] [B]^2} = 4.0 
2C + 5D → 2A + 4B …… K₂ = [A]^2 [B]^4 / {[C]^2 [D]^5} = ? 

K₁ = 4.0 
[C] [D]^(5/2) / {[A] [B]^2} = 4.0 

Reciprocal of the above expression : 
[A] [B]^2 / {[C] [D]^(5/2)} = 1 / 4.0 
[A] [B]^2 / {[C] [D]^(5/2)} = 0.25 

Square of the above expression : 
[A]^2 [B]^4 / {[C]^2 [D]^5} = 0.25^2 
[A]^2 [B]^4 / {[C]^2 [D]^5} = 0.0625 
[A]^2 [B]^4 / {[C]^2 [D]^5} ≈ 0.063 (to 2 sig. fig.) 

Hence, K₂ = [A]^2 [B]^4 / {[C]^2 [D]^5} = 0.063